OpenStudy (anonymous):

a^3+3a^2=3a^2+3b^2+3c^2-25 b^3+3b^2=3a^2+3b^2+3c^2-25 c^3+3c^2=3a^2+3b^2+3c^2-25 what is abc?

geerky42 (geerky42):

I think I saw this problem on Brilliant.org

OpenStudy (anonymous):

oh really?

geerky42 (geerky42):

Yeah... look so familiar...

OpenStudy (anonymous):

how can i find a+b+c?

OpenStudy (anonymous):

it's only that i am having problem with

geerky42 (geerky42):

never mind, I think of this: https://brilliant.org/community-problem/makes-me-wonder/?group=vRrE7dHGIKhr

geerky42 (geerky42):

Just clear things up... \[ \Large a^3+3a^2=3a^2+3b^2+3c^2-25 \\ \Large b^3+3b^2=3a^2+3b^2+3c^2-25 \\ \Large c^3+3c^2=3a^2+3b^2+3c^2-25\]

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

\[a +b +c =-3\] and I don't get why

geerky42 (geerky42):

oh you don't need to find abc?

OpenStudy (anonymous):

well i do need that but i have the solution to this problem and this is the only part that i don't understand

geerky42 (geerky42):

@mathslover @e.mccormick @ganeshie8 @ParthKohli Can you help?

OpenStudy (anonymous):

can you guys plz help?

OpenStudy (anonymous):

i dont really know. @hoblos @e.mccormick @whpalmer4 can you help him?

mathslover (mathslover):

Well, I am trying to find ... Please give me some time.